User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!


Hobart 26m

Module: HOB+0132, start pos ? GB

  • 00:00 UT - Experiment started OK
  • 12:18 UT - Lost connection to drive pc, restarted both drive pc and field system pc, but no luck. Called Jamie, he is trying to fix it. Probably related to the network still not being quite right. Brett restarted some switches out at Mt Pleasant, but issues still remain. (Patrick)
  • 13:44 UT - Restarted schedule. Problem was a struggling DNS server that was responding to queries slowly. Jamie pointed Newsmerd and Hobart to use a different DNS server which fixed the issues. Resumed with scan 344-1344. Missed scans were 344-1218a - 344-1341 (Patrick)
  • 14:51 UT - Windstow until 15:19. Had to send source=stow command again, as telescope was not stowing. (Patrick)
  • 15:57 UT - Windstow until 16:59. (Patrick)
  • 17:48 UT - Windstow until 17:59. (Patrick)
  • 23:06 UT - Schedule halted for cryo maintenance. (JMc)
  • 23:36 UT - Schedule resumes. All scans between 344-2305b and 344-2336 invalid (JMc).


Module: GSFC+024, start pos 0 GB

  • 00:00 UT - Experiment started OK
  • 00:01 UT - First scan (344-0001) recorded onto previous experiment's module (GSFC+012). Scans from 344-0003 onward recorded onto correct module. This means that will lie about data difference (Patrick)
  • 13:00 UT - UTAS VPN to Ke back online, all seems to be good. (Patrick)


Module: GSFC+023, start pos 0 GB

  • 00:00 UT - Experiment started OK
  • 12:15 UT - UTAS VPN to Yg back online, all seems to be good. (Patrick)
/home/www/auscope/opswiki/data/pages/handover/c1713_v1713.txt · Last modified: 2017/12/10 23:35 by Jamie McCallum